# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ III.4 Slow quenches and Kibble-Zurek mechanism
These experiments suggest that in the regime where the quench dynamics is consistent with the KZ theory, the scaling is described by the mean-field critical model. Indeed, the experiments are performed with laser cooled ions at Doppler temperatures at which quantum effects are not expected to be visible (see Sec. IIA and Ref. [294]). Quantum effects could be measured for slower quenches and a purely coherent dynamics. In this case it should be possible to observe the scaling of defects determined by the critical exponents of the Ising model [296]. In Ref. [297] a DMRG simulation reported the crossover from classical to quantum KZ scaling, where the scaling exponents are the critical exponent of the quantum critical model, identifying the requirements on the quench rate.
@@ III.5 Stick-slip motion and nano-friction
The possibility to image the ions within a Coulomb crystal and to control the forces at play make these systems ideal platforms for investigating static friction. Static friction refers to the finite work needed to put in motion an object in contact with a surface. It is associated with dynamical phenomena such as stick-slip motion, which consists of intermittent sliding and stuck phases and can characterize the motion of a mesoscopic object dragged over a surface. Stick-slip motion has been observed in laser–cooled ion crystals rotating in Penning traps in a setup shown in Fig. 47. Stress was created by two opposing torques generated by a rotating electric field that controls the ion crystal rotation through forces applied to the crystal boundary, and radiation pressure forces from a focused laser beam, implementing Doppler cooling. Stick-slip dynamics was manifested in sudden angular jumps (“slips”) of the crystal orientation alternating with intervals when the crystal orientation was phase locked (“stuck”) relative to the rotating wall perturbation, as seen in Fig. 48. Interestingly, the slip amplitudes were shown to exhibit a power-law distribution reminiscent of self-organized critical phenomena.
@@
Figure 48: Crystal orientation in the frame of the rotating perturbation for measurements with two different perpendicular beam torques. The torque is greater in measurement 2. The lines are from a linear regression fit. The inset shows a magnified plot of the first 60 s of data. From 216.
@@
Figure 49: The probability of scattering a photon from a one-dimensional optical lattice as a function of the optical lattice depth for a small one- (top), two- (center), and a three-dimensional Coulomb crystal (bottom). The red and blue triangles are experimental data points for red- and blue-detuned lattices, respectively. As expected for pinned ions, the scattering probability is higher for pinning at intensity maxima (red detuning) than at intensity minima (blue detuning). For comparison, the red and blue dashed lines show the theoretical scattering probabilities expected for delocalized ions. From 188, which includes a discussion of the error bars and uncertainties.
@@
Figure 50: Ion-crystal emulator of stick-slip friction. (a) A Coulomb crystal of 174Yb+ ions subjected to an optical lattice potential. Ions are imaged with a single-ion-resolving microscope. The typical ion spacing is 6 $\mu m$, and the lattice period is aa=185 nm. In the bottom illustration of the corrugated potential, the lattice period and the corrugation are strongly exaggerated. (b) Stick-slip results from bistability, illustrated here for a single ion. A shear force is linearly ramped, causing the ion to jump between minima. The stages of the stick-slip process are labeled by numbered diamonds and illustrate the dynamics of the ion during the ramp. The ion’s position is extracted from its fluorescence, which is proportional to the lattice potential energy. (c) Fluorescence versus applied force during the forward transport (green squares) and reverse transport (red circles), showing hysteresis that is used to measure the maximum static friction force $F_{s}$. The bold data points indicate the ion’s position before a slip, and only those data are used to reconstruct the force-displacement curve. (d) The force-displacement hysteresis loop encloses an area equal to twice the dissipated energy per slip $\Delta W$. The unit $m\nu^{2}a$ of the applied force corresponds to $2.8\times 10^{-19}$ N; here, $\nu=2\pi\times 364$ kHz. Adapted from 44.
@@
Stick-slip motion has also been observed in ion chains by realizing a more controlled environment. Pinning at the sub-wavelength scale can be observed using resonance fluorescence and has been measured in small ion clusters confined by optical lattices [197, 104, 188, 43, 115]. Figure 49 shows the resonance fluorescence of trapped ions as a function of lattice depth for various geometries, allowing the identification of the ions’ positions within the lattice. The dragging of an object over a surface can be implemented by setting the chain in motion relative to the optical lattice, as illustrated in Fig. 50: Stick-slip motion results in bistability in the potential energy. Bistability manifests itself as hysteresis in the ion position, from which the maximum static friction force can be extracted. In Ref. [44] the friction force was measured across five orders of magnitude in velocity by changing the depth of the lattice, demonstrating the versatility of ion chains as platforms for studying nanofriction. A further advantage of these realizations is that they are amenable to detailed modeling [117, 256, 18, 63, 325]. In particular, when the ion chain can be described by an elastic crystal, the dynamics is captured by a paradigmatic model of static friction, the so-called Frenkel-Kontorova (FK) model [33, 256, 63]:
$V_{\rm FK}=\frac{1}{2}\sum_{i,j}{\mathcal{K}}_{ij}(q_{i}-q_{j})^{2}+\frac{U}{2}(1-\cos(2\pi x_{j}/a))\,$, (68)
@@
For uniform chains and nearest-neighbor couplings, when $x_{j}^{(0)}=jd_{0}$ and ${\mathcal{K}}_{ij}=K_{0}\delta_{j,i+1}$, the ground state and dynamical properties of the FK model are fully determined by two dimensionless parameters, the ratio $U/K_{0}$ between the substrate potential and the elastic force, and the ratio $d_{0}/a$ between the two characteristic length scales. At fixed, incommensurable values of the ratio $d_{0}/a$, the critical value $U/K_{0}]_{c}$ signals the Aubry transition, separating the sliding phase, where forces giving rise to sticking cancel, from the pinned phase, where the motion alternates stick and slip phases and the minimal force needed to initiate sliding is finite [33]. In the pinned phase, continuous translational invariance is broken and pinning is signaled by a nonzero frequency of the lowest phonon mode. At fixed lattice depth, a given commensurate (periodic) phase at a commensurable ratio $d_{0}/a$ remains stable for small variations $d/a=d_{0}/a\pm\delta$. The system undergoes a discontinuous transition to an incommensurate (sliding) phase at a critical value $\delta_{c}>0$ of the deviation from commensurability, where formation of solitons is energetically favorable and the phase becomes sliding.
@@
Figure 51: Commensurate-incommensurate phase transition in an extension of the setup of [43], for two (left) and five ions (right). The plots are the measured kink density $p_{k}$ (top), deformation $|p_{0}-p_{k}|$ (middle), and static friction force $F_{\mathrm{s}}$ (bottom), vs mispit $p_{0}$. Lines in (a) and (b) are the kink density $p_{k}=p_{0}$ for $U=0$. Solid curves in (a)-(f) are numerical simulations for parameters matching those of the data shown, using $k_{\mathrm{B}}T/U=0.2$ [(a), (c), and (e)] and 0.5 [(b), (d), and (f)]. Dashed curves are $T=0$ simulations. $F_{\mathrm{max}}$ is the maximal friction in the single-slip regime. From 115.
Note that in the Paul trap the friction force does not vanish in the sliding phase, but becomes independent of the mismatch $p_{0}$ or of the lattice depth $U$. In this environment, the sliding-pinning Aubry transition breaks the reflection symmetry about the trap center [31, 18, 110].
The Aubry transition was observed in the setup of Fig. 50 by measuring the static friction force as a function of optical lattice depth [43]. It was also reported in a setup [172], which simulated the FK model in a chain of about 30 ions in the absence of a periodic lattice. In the realisation of [172], in fact, the elastic crystal and the substrate are the upper and the lower chain of an extended kink, which is metastable in the zigzag phase for a finite range of aspect ratios, see Fig. 52(a). In this setup, a cooling laser induces differential forces between the upper and the lower chain and can thus set one into motion with respect to the other [172, 173]. The parallel chains forming the extended kink can slide on top of one another. By decreasing the transverse trap frequency, the extended kink becomes unstable and the stable structure is a zigzag, corresponding to locking one chain with respect to the other, realizing a pinned phase. The symmetry breaking is revealed by measuring the relative axial distance $\Phi$ between the ions of the different layers as a function of the trap aspect ratio. The results of the measurement are shown in Fig. 52(b). Panel (c) reports the measurement of the frequency of the lowest vibrational mode, which is detected by a second laser performing vibrational spectroscopy. The smoothing of the transition is attributed to thermal fluctuations and is reproduced by both molecular dynamics simulations and a microscopic model that incorporates thermal effects [174].
@@
Despite the analogy with the FK model, the two chains in Fig. 52 mutually interact and hence proper modeling of their dynamics must take into account the back action of the elastic crystal on the surface on which it is dragged. This is indeed closer to the typical scenario of nanofriction than the idealized FK model. Another implementation of deformable substrates is realized when the ion chains interact with the standing wave field of an optical cavity [130, 169, 50, 49, 197]. For sufficiently large couplings, the cavity optical lattice depth depends nonlinearly on the ion density, such that it is constant in the sliding phase and can become steeper (or shallower, depending on the experimental parameters) in the pinned phase [62, 63, 110]. For this setup the sliding and pinned phases are separated by bistable regions where superlubric and stick-slip dynamics coexist [110]. Moreover, photon-scattering into the cavity mode can cool the chain vibrations even down to the zero-point motion. Entanglement in the pinned phase has been theoretically studied in [165]. Phase, temperatures, and correlations can be measured in the spectrum of the intensity at the cavity output [63, 112].
@@ III.6 Kink dynamics and quantum sliding
The level of control suggests that it could be possible to observe quantum dynamics of kinks. This question is at the center of studies in material science of quantum lubrication, where sliding is initiated by quantum tunnelling [179, 325]. Path Integral Quantum Monte Carlo simulations identified the regime of parameters where quantum-induced sliding phases could be observed in chains of 5 ions [30]. Semiclassical analysis discussed the dynamics of tunnelling of the kink in the platform of [172], see [314, 316].
A general framework for quantum lubrication in ion chains was developed in [186, 208, 209], which permit to derive a quantum field theory for the solitons as quasiparticles. Differing from the short-range Frenkel-Kontorova model, the long-range Coulomb interactions give rise to important features that substantially modify the ground state of the chain, and thus the Aubry and the commensurate-incommensurate transition. The first step consists in the determination of the functional form of the topological kink. In the limit of large chains the ions interparticle distance $d_{0}$ can be assumed to be uniform and the kink is described by a finite displacement $\delta x_{j}$ from a commensurate configuration, that is conveniently cast in terms of a phase $\theta_{j}=2\pi\delta x_{j}/a$, see Fig. 53(a) and (b). For $d_{0}\gg a$, taking the continuum limit gives the field $\theta(x,t)$ as a function of the position $x$ along the chain, which is the solution of the modified sine-Gordon equation [32, 186]:
@@
and where $x$ is here rescaled by $a$. The solutions of Eq. (70) are constrained to satisfy the mismatch $\delta=\min_{n_{0}\in\mathbb{N}}(d_{0}-n_{0}a)/a$. The equation is parametrized by the velocity $v_{s}=\sqrt{{3K_{0}}/({2m})}$ and the mass term $M=\sqrt{{8\pi^{2}U}/({3K_{0}a^{2}})}$, with $K_{0}=2Q^{2}/(4\pi\epsilon_{0}d_{0}^{3})$. In the short-range FK model the integral term vanishes and the solution is a sine-Gordon soliton with sound velocity $v_{s}$ and length proportional to $d_{0}/M$, while the mass $M$ scales the energy of a kink. The integral term in Eq. (70) is due to the long-range coupling and predicts that the soliton decays with a power-law tail. As a consequence, the effective width of the sine-Gordon soliton scales with the size of the chain as $\log N$ [186]. Moreover, solitons interact with power-law, Coulomb interactions [208]. This property changes the nature of the commensurate-incommensurate phase transition with respect to the short-range model, as shown in [208, 209].
In [208] a quantum field theoretical model was developed for the Coulomb solitons $\theta(x,t)$, which permits to map the ground state of the Wigner chain in the periodic potential to a massive Thirring model, where the solitons are charged fermionic excitations over an effective Dirac sea, as illustrated in Fig. 53(c). The Thirring Hamiltonian, in turn, is equivalent to a spin model, more precisely, to a XXZ Hamiltonian of spin-$\frac{1}{2}$ particles in an external field, where the solitons are long-range interacting defects, as shown in Fig. 53(d).
Notably, in this theory quantum effects in the soliton’s dynamics scale with the effective Planck constant $\beta^{2}$ [208],
@@
which is proportional to the squared root of the ratio between kinetic and Coulomb energy. This quantity is proportional to the parameter $\tilde{\hbar}$ introduced in Sec. II.1.5, and is scaled by the ratio $(d_{0}/a)^{2}$ between the characteristic length scales of chain and lattice. Therefore, quantum effects can be tuned by changing the ratio $d_{0}/a$, and can take values ranging from $\beta^{2}\sim 10^{-4}$, deep in the mean-field regime, to $\beta^{2}\sim 0.1$, where sliding induced by quantum tunneling could be measured.
Figure 53: (a) Schematic representation of the Wigner crystal confined by an optical lattice. The substrate potential leads to dislocation of the ions from the crystal’s equilibrium position. (b) The dislocations (solitons) are displayed as phase shift $\theta(x)$ along the chain axis $x$ (orange). The first-order derivative of $\theta(x)$ (blue) displays local maxima where the solitons are located. The yellow drawing illustrates that the solitons behave as interacting charges. (c) In the Thirring model the solitons are positive-energy excitations over a filled Dirac sea, where the gap is twice the soliton mass $M_{0}$. The phase is commensurate when the chemical potential $h^{\prime}$ (controlled by the misfit) falls within the gap. For $h^{\prime}>M_{0}$, the solitons proliferate and the phase is incommensurate. (d) Interpretation of the spin-fermion equivalence in the discretized Thirring model, corresponding to a long-range interacting XXZ antiferromagnetic spin chain: Solitons are defects (in orange) in the staggered ordering of a spin chain. From 208.
A mean-field study permits to determine analytically the regions of stability of the commensurate configurations at fixed $N$ as a function of the mismatch $\delta$ and of the temperature $T$ [209]. Using the parameters of Ref. [43], for a chain of 100 ions 173Yb+ with interparticle distance $d_{0}=6~\mu\mathrm{m}$ and lattice periodicity $a=185~\mathrm{nm}$, commensurate phases can be observed at temperatures $T\lesssim 1\,\mathrm{mK}$ [209]. Quantum effects could be revealed by means of spectroscopic measurements for relative precisions below the effective Planck’s constant $\beta^{2}$.

### v2

@@ III.4 Slow quenches and Kibble-Zurek mechanism
The agreement between experiments and molecular dynamics simulations shows that the observed scalings emerge from classical dynamics. Indeed, the experiments are performed with laser-cooled ions at Doppler temperatures at which quantum effects are not expected to be visible (see Sec. IIA and Ref. [328]). Quantum effects could be measured for slower quenches and purely coherent dynamics. In this case, it should be possible to observe the scaling of defects determined by the critical exponents of the Ising model [331], namely, the quantum KZ mechanism. In Ref. [332] a DMRG simulation was used to determine the defect statistics generated by slow quenches across the quantum critical model of Sec. II.1.5. A crossover from classical to quantum KZ scaling was reported as a function of the quench rate, identifying the requirements on the ramp speed in order to unveil quantum effects on defect formation.
@@ III.5 Stick-slip motion and nano-friction
The possibility to image the ions within a Coulomb crystal and to control the forces at play make these systems ideal platforms for investigating static friction [360]. Static friction refers to the finite work needed to put in motion an object in contact with a surface. It is associated with dynamical phenomena such as stick-slip motion, which consists of intermittent sliding and stuck phases and can characterize the motion of a mesoscopic object dragged over a surface. Stick-slip motion has been observed in laser–cooled ion crystals rotating in Penning traps in the setup shown in Fig. 47. Stress was created by two opposing torques generated by a rotating electric field that controls the ion crystal rotation through forces applied to the crystal boundary, and radiation pressure forces from a focused laser beam, implementing Doppler cooling [240]. Stick-slip dynamics was manifested in sudden angular jumps (“slips”) of the crystal orientation alternating with intervals when the crystal orientation was phase locked (“stuck”) relative to the rotating wall perturbation, as seen in Fig. 48. Interestingly, the slip amplitudes were shown to exhibit a power-law distribution reminiscent of self-organized critical phenomena [240].
@@
Figure 48: Crystal orientation in the frame of the rotating perturbation for measurements with two different perpendicular beam torques. The torque is greater in measurement 2. The lines are from a linear regression fit. The inset shows a magnified plot of the first 60 s of data. From 240.
@@
Figure 49: (Color online) The probability of scattering a photon from a one-dimensional optical lattice as a function of the optical lattice depth for a small one- (top), two- (center), and a three-dimensional Coulomb crystal (bottom). The red (up-pointing) and blue (down-pointing) triangles are experimental data points for red- and blue-detuned lattices, respectively. As expected for pinned ions, the scattering probability is higher for pinning at intensity maxima (red detuning) than at intensity minima (blue detuning). For comparison, the red (lower) and blue (upper) dashed lines show the theoretical scattering probabilities expected for delocalized ions. From 205, which includes a discussion of the error bars and uncertainties.
@@
Figure 50: (Color online) Ion-crystal emulator of stick-slip friction. (a) A Coulomb crystal of 174Yb+ ions subjected to an optical lattice potential. Ions are imaged with a single-ion-resolving microscope. The typical ion spacing is 6 $\mu m$, and the lattice period is aa=185 nm. In the bottom illustration of the corrugated potential, the lattice period and the corrugation are strongly exaggerated. (b) Stick-slip results from bistability, illustrated here for a single ion. A shear force is linearly ramped, causing the ion to jump between minima. The stages of the stick-slip process are labeled by numbered diamonds and illustrate the dynamics of the ion during the ramp. The ion’s position is extracted from its fluorescence, which is proportional to the lattice potential energy. (c) Fluorescence versus applied force during the forward transport (green squares) and reverse transport (red circles), showing hysteresis that is used to measure the maximum static friction force $F_{s}$. The bold data points indicate the ion’s position before a slip, and only those data are used to reconstruct the force-displacement curve. (d) The force-displacement hysteresis loop encloses an area equal to twice the dissipated energy per slip $\Delta W$. The unit $m\nu^{2}a$ of the applied force corresponds to $2.8\times 10^{-19}$ N; here, $\nu=2\pi\times 364$ kHz. Adapted from 47.
@@
Ion chains offer a controlled environment for studying static friction. Pinning at the sub-wavelength scale can be observed using resonance fluorescence and has been measured in small ion clusters confined by optical lattices [111, 216, 205, 46, 125]. Figure 49 shows the resonance fluorescence of trapped ions as a function of lattice depth for various geometries, allowing the identification of the ions’ positions within the lattice. The dragging of an object over a surface can be implemented by setting the chain in motion relative to the optical lattice, as illustrated in Fig. 50: Stick-slip motion results in bistability in the potential energy and manifests itself as hysteresis in the ion position. In Ref. [47, 124] the maximal friction force was extracted from measurements of the hysteresis of the ion position.
A further advantage of studying nano-friction with ion chains is that these platforms are amenable to microscopic modeling and mapping to paradigmatic models [127, 21, 287, 68, 360]. In particular, when the ion chain can be described by an elastic crystal, the dynamics is captured by a paradigmatic model of static friction, the so-called Frenkel-Kontorova (FK) model [35, 287, 68]:
$V_{\rm FK}=\frac{1}{2}\sum_{i,j}{\mathcal{K}}_{ij}(q_{i}-q_{j})^{2}+\frac{U}{2}\sum_{j}(1-\cos(2\pi x_{j}/a))\,$, (68)
@@
For uniform chains and nearest-neighbor couplings, when $x_{j}^{(0)}=jd_{0}$ and ${\mathcal{K}}_{ij}=K_{0}\delta_{j,i+1}$, the ground state and dynamical properties of the FK model are fully determined by two dimensionless parameters, the ratio $U/K_{0}$ between the substrate potential and the elastic force, and the ratio $d_{0}/a$ between the two characteristic length scales. At fixed, incommensurable values of the ratio $d_{0}/a$, the critical value $U/K_{0}]_{c}$ signals the Aubry transition, separating the sliding phase, where forces giving rise to sticking cancel, from the pinned phase, where the motion alternates stick and slip phases and the minimal force needed to initiate sliding is finite [35]. In the pinned phase, continuous translational invariance is broken and pinning is signaled by a nonzero frequency of the lowest phonon mode. At fixed lattice depth, a given commensurate (periodic) phase at a commensurable ratio $d_{0}/a$ remains stable for small variations $d/a=d_{0}/a\pm\delta$, where the mismatch $\delta$ quantifies the deviation from the commensurable ratio. The system undergoes a discontinuous transition to an incommensurate (sliding) phase at a critical value $\delta_{c}>0$ where formation of solitons is energetically favorable and the phase becomes sliding.
@@
Figure 51: Commensurate-incommensurate phase transition in an extension of the setup of [46], for two (left) and five ions (right). The plots are the measured kink density $p_{k}$ (top), deformation $|p_{0}-p_{k}|$ (middle), and static friction force $F_{\mathrm{s}}$ (bottom), vs misfit $p_{0}$. Lines in (a) and (b) are the kink density $p_{k}=p_{0}$ for $U=0$. Solid curves in (a)-(f) are numerical simulations for parameters matching those of the data shown, using $k_{\mathrm{B}}T/U=0.2$ [(a), (c), and (e)] and 0.5 [(b), (d), and (f)]. Dashed curves are $T=0$ simulations. $F_{\mathrm{max}}$ is the maximal friction in the single-slip regime. From 125.
Note that in the Paul trap the friction force does not vanish in the sliding phase, but becomes independent of the misfit $p_{0}$ or of the lattice depth $U$. In this environment, the sliding-pinning Aubry transition breaks the reflection symmetry about the trap center [34, 21, 119].
The Aubry transition was observed in the setup of Fig. 50 by measuring the static friction force as a function of optical lattice depth [46]. It was also reported in a setup [187], which simulated the FK model in a chain of about 30 ions in the absence of a periodic lattice. In the realization of [187], in fact, the elastic crystal and the substrate are the upper and the lower chain of an extended kink, which is metastable in the zigzag phase for a finite range of aspect ratios, see Fig. 52(a). In this setup, a cooling laser induces differential forces between the upper and the lower chain and can thus set one into motion with respect to the other [187, 188]. The parallel chains forming the extended kink can slide on top of one another. By decreasing the transverse trap frequency, the extended kink becomes unstable and the stable structure is a zigzag, corresponding to locking one chain with respect to the other, realizing a pinned phase. The symmetry breaking is revealed by measuring the relative axial distance $\Phi$ between the ions of the different layers as a function of the trap aspect ratio. The results of the measurement are shown in Fig. 52(b). Panel (c) reports the measurement of the frequency of the lowest vibrational mode, which is detected by a second laser performing vibrational spectroscopy. The smoothing of the transition is attributed to thermal fluctuations and is reproduced by both molecular dynamics simulations and a microscopic model that incorporates thermal effects [189].
@@
Despite the analogy with the FK model, the two chains in Fig. 52 mutually interact and hence proper modeling of their dynamics must take into account the back action of the elastic crystal on the surface on which it is dragged. This is indeed closer to the typical scenario of nanofriction, where the substrate is deformable, than the idealized FK model. Another implementation of deformable substrates is realized when the ion chains interact with the standing wave field of an optical cavity [141, 184, 216, 54, 53]. For sufficiently large couplings, the cavity optical lattice depth depends nonlinearly on the ion density, such that it is constant in the sliding phase and can become steeper (or shallower, depending on the experimental parameters) in the pinned phase [67, 68, 119]. For this setup the sliding and pinned phases are separated by bistable regions where superlubric and stick-slip dynamics coexist [119]. Moreover, photon-scattering into the cavity mode can cool the chain vibrations even down to the zero-point motion [121]. Entanglement in the pinned phase has been theoretically studied in [178]. Phase, temperatures, and correlations can be measured in the spectrum of the intensity at the cavity output [68, 121].
@@ III.6 Kink dynamics and quantum sliding
The experimental level of control suggests that it could be possible to observe quantum dynamics of kinks. This question is at the center of studies in material science of quantum lubrication, where sliding is initiated by quantum tunnelling [195, 360]. Path Integral Quantum Monte Carlo simulations identified the regime of parameters where quantum-induced sliding phases could be observed in chains of 5 ions [33]. Semiclassical analysis discussed the dynamics of tunnelling of the kink in the platform of [187], see [349, 351].
A general framework for quantum lubrication in ion chains was developed in [197, 232, 233], which permits one to derive a quantum field theory for the solitons as quasiparticles. These works showed that the long-range Coulomb interactions substantially modify the features of the topological kinks with respect to the short-range FK model, favoring soliton formation in long chains [232, 233].
Figure 53(a) sketches a chain displaying two kinks. The kink configuration is described by the finite displacement $\delta x_{j}$ from the position of the commensurate configuration and can be cast in terms of a phase $\theta_{j}=2\pi\delta x_{j}/a$, such that a change by $2\pi$ corresponds to a kink, see Fig. 53(b) and [59]. For long chains and $d_{0}\gg a$ the phase $\theta_{j}$ is captured by the field $\theta(x,t)$, which is the solution of the modified sine-Gordon equation [36, 197]:
@@
and where $x$ is here rescaled by $a$. The equation is parametrized by the velocity $v_{s}=\sqrt{{3K_{0}}/({2m})}$ and the mass term $M=\sqrt{{8\pi^{2}U}/({3K_{0}a^{2}})}$, with $K_{0}=2Q^{2}/(4\pi\epsilon_{0}d_{0}^{3})$. The solutions of Eq. (70) are constrained to satisfy the mismatch $\delta=\min_{n_{0}\in\mathbb{N}}(d_{0}-n_{0}a)/a$. In the short-range FK model the integral term vanishes and the solution is a sine-Gordon soliton with sound velocity $v_{s}$ and length proportional to $d_{0}/M$, while the mass $M$ scales the energy of a kink. The Coulomb interactions give rise to the integral term. As a consequence, at the center the kink has the form of a sine-Gordon soliton whose width now scales with the size of the chain as $\log N$ while it asymptotically decays as $1/r$ with the distance $r$ from its center [197]. This asymptotic decay gives rise to Coulomb-like interactions between distant solitons [232].
Equation (70) permits to map the dynamics of a frustrated ion chain to a quantum field theoretical model, the massive Thirring model, where the commensurate phase is an effective Dirac sea and the solitons are charged fermionic excitations, as illustrated in Fig. 53(c). The chemical potential is proportional to the mismatch and shows that, when this exceeds the gap, excitations (solitons) are generated. Because of the Coulomb interactions, the chemical potential also depends on the chain’s size: $h^{\prime}\propto\delta\ln N$. This scaling of $h^{\prime}$ indicates that the Coulomb interactions dictate the order in the thermodynamic limit. Correspondingly, the critical mismatch $\delta_{c}$ shrinks as $1/\ln N$, giving rise to interaction-induced sliding [232, 233]. An equivalent spin model can be derived from the Thirring Hamiltonian, which is amenable to numerical analyses. This is a long-range XXZ Hamiltonian of spin-$\frac{1}{2}$ particles in an external field. The solitons are mapped into long-range interacting defects, as shown in Fig. 53(d). In these models quantum effects in the soliton’s dynamics scale with the effective Planck constant $\beta^{2}$ [232],
@@
which is proportional to the square root of the ratio between kinetic and Coulomb energy. This quantity is proportional to $\tilde{\hbar}$ introduced in Sec. II.1.5, which determines the size of quantum effects at the linear-zigzag transition. The proportionality factor contains the ratio $(d_{0}/a)^{2}$ between the characteristic length scales of chain and lattice and suggests that size of quantum effects can be tuned by changing the trap aspect ratio and/or the wavelength of the confining lattice. Taking experimentally accessible parameters, it can range from $\beta^{2}\sim 10^{-4}$, deep in the mean-field regime, to $\beta^{2}\sim 0.1$, where quantum sliding could be measured [232, 233].
Figure 53: (Color online) (a) Schematic representation of the Wigner crystal in the presence of an externally applied optical lattice, mimicking a substrate potential. The substrate potential leads to dislocations of the ions from the crystal’s equilibrium position. (b) The dislocations (solitons) are displayed as phase shift $\theta(x)$ along the chain axis $x$ (monotonous orange curve). The first-order derivative of $\theta(x)$ (blue) exhibits local maxima where the solitons are located. The yellow dashed curves illustrate that the solitons behave as interacting charges. (c) In the Thirring model the solitons are positive-energy excitations over a filled Dirac sea, where the gap is twice the soliton mass $M_{0}$. The phase is commensurate when the chemical potential $h^{\prime}$ (controlled by the mismatch) falls within the gap. For $h^{\prime}>M_{0}$, the solitons proliferate and the phase is incommensurate. (d) Interpretation of the spin-fermion equivalence in the discretized Thirring model, corresponding to a long-range interacting XXZ antiferromagnetic spin chain: Solitons are defects (co-polarized adjacent spins highlighted in orange) in the staggered ordering of a spin chain. From 232.
A mean-field study predicted the regions of stability of the commensurate configurations at fixed $N$ as a function of the mismatch $\delta$ and of the temperature $T$ [233]. Using the parameters of Ref. [46], for a chain of 100 ions 173Yb+ with interparticle distance $d_{0}=6~\mu\mathrm{m}$ and lattice periodicity $a=185~\mathrm{nm}$, commensurate phases can be observed at temperatures $T\lesssim 1\,\mathrm{mK}$ [233]. Quantum sliding could be revealed by means of spectroscopic measurements for relative precisions below the effective Planck’s constant $\beta^{2}$.